# Vector Quantization using Gaussian Variational Autoencoder

### v1

@@ Abstract
Vector quantized variational autoencoder (VQ-VAE) is a discrete auto-encoder that compresses images into discrete tokens. It is difficult to train due to discretization. In this paper, we propose a simple yet effective technique, dubbed Gaussian Quant (GQ), that converts a Gaussian VAE with certain constraint into a VQ-VAE without training. GQ generates random Gaussian noise as a codebook and finds the closest noise to the posterior mean. Theoretically, we prove that when the logarithm of the codebook size exceeds the bits-back coding rate of the Gaussian VAE, a small quantization error is guaranteed. Practically, we propose a heuristic to train Gaussian VAE for effective GQ, named target divergence constraint (TDC). Empirically, we show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures. Furthermore, TDC also improves upon previous Gaussian VAE discretization methods, such as TokenBridge. The source code is provided in https://github.com/tongdaxu/VQ-VAE-from-Gaussian-VAE.
@@ 1 Introduction
Vector-quantized variational autoencoder (van2017neural) is an autoencoder that compresses images into discrete tokens. It is fundamental to autoregressive generative models (esser2021taming; chang2022maskgit; yu2023language; sun2024autoregressive). However, VQ-VAE is difficult to train: the encoding process of VQ-VAE is not differentiable and challenges such as codebook collapse often emerge (sonderby2017continuous). Special techniques are required to ensure the convergence of VQ-VAE, such as commitment loss (van2017neural), expectation maximization (EM) (roy2018theory), Gumbel-Softmax (jang2016categorical; maddison2016concrete; sonderby2017continuous), and entropy loss (yu2023language; zhao2024image).
In this paper, we circumvent the challenge of training VQ-VAE by converting a Gaussian VAE with certain constraint into a VQ-VAE without any training. More specifically, we propose Gaussian Quant (GQ), a simple yet effective method for training-free conversion. The core idea is to generate a codebook of one-dimensional Gaussian noise and, for each dimension of the posterior, select the codebook entry that is closest to the posterior mean. Theoretically, we show that as the logarithm of the codebook size exceeds the bits-back coding bitrate (hinton1993keeping; townsend2019practical) of the Gaussian VAE, the resulting quantization error is small. In other words, GQ and the Gaussian VAE exhibit similar rate-distortion performance. This result serves as the theoretical foundation of GQ and provides a principled guideline for selecting codebook sizes.
Practically, we introduce the target divergence constraint (TDC) to train a Gaussian VAE for efficient conversion. TDC encourages the Gaussian VAE to achieve the same Kullback–Leibler (KL) divergence for each dimension, corresponding to the bits-back coding bitrate. Empirically, we demonstrate that GQ with Gaussian VAE trained by TDC, outperforms previous VQ-VAEs such as VQGAN, FSQ, LFQ, and BSQ (van2017neural; mentzer2023finite; yu2023language; zhao2024image) in terms of reconstruction quality, using both UNet and ViT backbones. Additionally, we show that TDC can improve previous Gaussian VAE discretization methods, such as TokenBridge (wang2025bridging).
Our contributions can be summarized as follows:
  • (Section 3.1) We propose GQ, a simple yet effective approach that converts a pre-trained Gaussian VAE with certain constraint into VQ-VAE without training.
  • (Section 3.2) Theoretically, we prove that when the codebook size of GQ is close to the bits-back coding bitrate of the Gaussian VAE, the conversion error remains small.
  • (Section 3.3) Empirically, we introduce target divergence constraint (TDC) to implement GQ and show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures.
  • (Section 3.5) Furthermore, we show that TDC can be used to improve previous Gaussian VAE discretization approaches, such as TokenBridge.
@@ 2 Preliminaries
We denote the original image as $X$, the latent variable as $Z$, the encoder as $f(\cdot)$, and the decoder as $g(\cdot)$. We use $\log(\cdot)$ to denote the natural logarithm (base $e$) and KL divergence as $D_{KL}(\cdot||\cdot)$. Similarly, we use $\log_{2}(\cdot)$ to denote the logarithm with base $2$, and KL divergence as $D_{KL(2)}(\cdot||\cdot)$.
@@ 2.1 Vector Quantized Variational Autoencoder
VQ-VAE (van2017neural) transforms source image into a series of integer tokens, which can be decoded using a codebook and decoder. To facilitate auto-regressive generation, it typically involves a deterministic transformation and a shared codebook across different tokens. More specifically, VQ-VAE maintains a codebook $c_{1:K}$ with size $K$ and a bitrate of $\log K$. The encoding process of VQ-VAE involves finding the closest codeword $c_{j}$ in $c_{1:K}$ to the encoder output $f(x)_{i}$ for each latent dimension $i$. Denote distortion as $\Delta(\cdot,\cdot)$, the optimization target of VQ-VAE is the rate-distortion function weighted by Lagrangian multiplier $\lambda$:
$\mathcal{L}_{VQ}=\lambda\underbrace{\log K}_{\textrm{bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(\hat{z}))]}_{\textrm{distortion}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{Reg}},$
@@
and $\mathcal{L}_{Reg}$ are regularization for VQ-VAE to converge, such as a combination of commitment loss and codebook loss (van2017neural) and Gumbel Softmax loss (sonderby2017continuous).
2.2 Gaussian Variational Autoencoder and Bits-Back Coding
The Gaussian VAE is a special type of VAE (kingma2013auto) with a prior $\mathcal{N}(0,I)$ and a fully factorized Gaussian posterior $q(Z|X)$. The encoding process of a Gaussian VAE simply involves sampling the latent variable $z_{i}\sim q(Z_{i}|X)$ for each latent dimension $i$. Assuming $\log p(X|Z=z)\propto(1/\lambda)\Delta(X,g(z))$, then the negative evidence lower bound (ELBO) of the Gaussian VAE is equivalent to a rate-distortion function of a bits-back coding bitrate term and a distortion term:
@@ 2.2 Gaussian Variational Autoencoder and Bits-Back Coding
The bitrate of $z_{i}$ is the bits-back coding bitrate, defined as $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$ (hinton1993keeping; townsend2019practical). This is because, when compressing $X$ losslessly, one can communicate $z_{i}$ using $D_{KL}(q(Z|X)||\mathcal{N}(0,I))$ nats for arbitrary precision.
@@ 3 Gaussian Quant: Vector Quantization using Gaussian VAE
3.1 Direct Quantization of Gaussian VAE
We propose an extremely simple technique to obtain a VQ-VAE from a Gaussian VAE: we directly generate one-dimensional Gaussian noise as the codebook for VQ-VAE (van2017neural) and quantize the posterior mean $\mu_{i}$ of the Gaussian VAE independently for each dimension $i$. Because the codebook consists entirely of samples from a Gaussian distribution, we refer to our approach as Gaussian Quant (GQ). Specifically, we randomly generate $K$ codebook values $c_{1:K}\sim\mathcal{N}(0,1)$, which is the same for each dimension. Then, for each dimension $i$, we select the $c_{j}$ that is closest to the posterior mean $\mu_{i}$ and denote the quantized value as $\hat{z}_{i}$:
@@ 3.2 Theoretical Relationship between the Codebook Size and Quantization Error
Why GQ works and how to select $K$ are not straightforward questions. Theoretically, we show that GQ preserves the rate-distortion property of the Gaussian VAE: when the bitrate $\log K$ matches the bits-back coding bitrate of the Gaussian VAE, the quantization error is small. More specifically, we show that the probability of large quantization error decays doubly exponentially as the codebook bitrate $\log K$ exceeds the bits-back coding bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
Theorem 1.
@@
Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i}$ and $\sigma_{i}$, respectively. Assuming that the product and sum satisfy $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with respect to the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding bitrate. i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-e^{t}\underbrace{\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}}_{\textup{constant}})}.$ (4)
@@
Conversely, when the codebook bitrate $\log K$ is smaller than the bits-back coding bitrate, the probability of large quantization error increases exponentially toward $1$. More specifically, we show that the probability of large quantization error increases exponentially when the codebook bitrate $\log K$ is lower than the bits-back bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
Theorem 2.
@@
The probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with respect to the number of nats $t$ by which the codebook bitrate $\log K$ is lower than the bits-back coding bitrate, i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\geq 1-e^{-t}\underbrace{\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}}_{\textup{constant}}.$ (5)
@@
Theorems 1 and 2 provide a principled guideline for choosing $K$, such that $\log K$ should be close to the bits-back bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$. In practice, setting $\log_{2}K=\left\lceil D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\right\rfloor$ typically yields a small enough reconstruction error, where $\left\lceil\cdot\right\rfloor$ denotes the rounding operator. Using a larger $K$ does not provide additional benefits, while a smaller $K$ increases the error significantly.
@@
3.3 Practical Implementation with Target Divergence Constraint
There are two challenges to make GQ practical. The first challenge is, if we want to construct a VQ-VAE with a specific codebook size $K$, how can we train a Gaussian VAE with corresponding KL divergence? The second challenge is, for a vanilla Gaussian VAE trained to minimize the loss in Eq.2, the values of $D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))$ vary significantly across dimensions $i$. How to train a Gaussian VAE with KL divergence remains close to $\log K$ across each dimension?
@@ 3.3 Practical Implementation with Target Divergence Constraint
$\mathcal{L}_{TDC}=\sum_{i=1}^{d}\lambda_{i}D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+\Delta(X,g(z)),$
$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textup{ bits},\\ \lambda_{\textrm{mean}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\in[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha,\log_{2}K+\alpha}]\textup{ bits},\\ \lambda_{\textrm{max}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textup{ bits}.\end{array}\right$. (9)
To determine $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$, we extend the heuristic in MIRACLE (havasi2018minimal) and HiFiC (Mentzer2020HighFidelityGI). More specifically, we initialize $\lambda_{\textrm{min}}=\lambda_{\textrm{mean}}=\lambda_{\textrm{max}}=1$, and update them according to the following rule:
$\lambda_{\textrm{min}}=\lambda_{\textrm{min}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ min}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textrm{ else }\lambda_{\textrm{min}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{mean}}=\lambda_{\textrm{mean}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ mean}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>\log_{2}K\textrm{ else }\lambda_{\textrm{mean}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{max}}=\lambda_{\textrm{max}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ max}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textrm{ else }\lambda_{\textrm{max}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$ (10)
where $\beta$ is the hyper-parameter controlling update speed. To avoid numerical issue, we further clip $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$ into a range of $10^{-3},10^{3}]$ after each update. In practice, we use $\alpha=0.5,\beta=1.01$. In Appendix B, we propose an alternative implementation of TDC using Lambert W function, which is less effective for ViT models.
3.4 Grouping to Multiple Dimensions
The vanilla VQ-VAE has three key parameters: codebook size, codebook dimension and number of token. Some previous alternative to VQ-VAE, such as LFQ and BSQ (yu2023language; zhao2024image) limit the codebook dimension to 1. To support codebook dimension greater than 11, we can group $m$ tokens into a single large token with codebook size $\log_{2}K=\left\lceil\sum_{l=i}^{i+m}D_{KL}(q(Z_{l}|X)||\mathcal{N}(0,1))\right\rfloor$. There are three grouping strategies to achieve this: post-quantization (PQ), post-training (PT), and training-based (TR), with different trade-offs in flexibility and performance. We brief these strategies in main text and provide details in Appendix C.2.
PQ is the most flexible grouping strategy and can be applied after GQ. For PQ, since the posterior of the Gaussian VAE $q(Z|X)$ is a factorized Gaussian, the quantization is also independent across dimensions. This means that we can trivially combine $m$ tokens into a larger one, by treating each token as an integer in $K^{1/m}$-based number system and aggregate them, which is the same as other one dimensional VQ-VAEs (chang2022maskgit; mentzer2023finite; zhao2024image).
PT is less flexible than PQ, as it can only be applied before GQ and after training the Gaussian VAE. For PT, we can view the one-dimensional GQ in Eq. 3 as the maximum likelihood estimator of a one-dimensional Gaussian. This approach can be extended to an $m$-dimensional diagonal Gaussian distribution. Additionally, for low bitrate cases, we observe that $m$-dimensional PT leads to low codebook usage. This is because $|\mu_{i}|$ is bounded by $\sqrt{2D_{KL}}$. When the $D_{KL}$ is small, some vector in the codebook that is far from 0 is never used. To address this, we introduce a regularization term weighted by $\omega$ to improve codebook usage by encouraging the selection $c_{j}$ that is far from 0:

### v2

@@ Abstract
Vector-quantized variational autoencoders (VQ-VAEs) are discrete autoencoders that compress images into discrete tokens. However, they are difficult to train due to discretization. In this paper, we propose a simple yet effective technique dubbed Gaussian Quant (GQ), which first trains a Gaussian VAE under certain constraints and then converts it into a VQ-VAE without additional training. For conversion, GQ generates random Gaussian noise as a codebook and finds the closest noise vector to the posterior mean. Theoretically, we prove that when the logarithm of the codebook size exceeds the bits-back coding rate of the Gaussian VAE, a small quantization error is guaranteed. Practically, we propose a heuristic to train Gaussian VAEs for effective conversion, named the target divergence constraint (TDC). Empirically, we show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures. Furthermore, TDC also improves previous Gaussian VAE discretization methods, such as TokenBridge. The source code is provided in the supplementary materials.
Machine Learning, ICML
@@ 1 Introduction
Vector-quantized variational autoencoders (VQ-VAEs) (Van Den Oord et al., 2017) compress images into discrete tokens and are fundamental to autoregressive generative models (Esser et al., 2021; Chang et al., 2022; Yu et al., 2023; Sun et al., 2024a). However, VQ-VAEs are difficult to train: the encoding process is non-differentiable, and issues such as codebook collapse often arise (Sønderby et al., 2017). As a result, special techniques are required, including commitment loss (Van Den Oord et al., 2017), expectation maximization (EM) (Roy et al., 2018), Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2016; Sønderby et al., 2017), and entropy loss (Yu et al., 2023; Zhao et al., 2024).
Figure 2: Qualitative results on ImageNet dataset at 0.25 bits-per-pixel (bpp). GQ has the most visually pleasing reconstruction.
In this paper, we circumvent the difficulty of training VQ-VAEs by not training them at all. Specifically, we propose Gaussian Quant (GQ), which first trains a constrained Gaussian VAE and then converts it into a VQ-VAE without additional training. For conversion, GQ generates a codebook of one-dimensional Gaussian noise and, for each posterior dimension, select the entry closest to the posterior mean. We show theoretically that when the logarithm of the codebook size exceeds the bits-back coding rate (Hinton and Van Camp, 1993; Townsend et al., 2019) of the Gaussian VAE, the quantization error is small, providing a principled guideline for codebook size selection.
To train a Gaussian VAE for effective conversion, we introduce the target divergence constraint (TDC), which encourages the Gaussian VAE to achieve the same Kullback–Leibler (KL) divergence for each dimension, corresponding to the bits-back coding bitrate. Empirically, we demonstrate that GQ with Gaussian VAE trained by TDC, outperforms previous VQ-VAEs such as VQGAN, FSQ, LFQ, and BSQ (Van Den Oord et al., 2017; Mentzer et al., 2023; Yu et al., 2023; Zhao et al., 2024) in terms of reconstruction, on UNet and ViT backbones. Additionally, we show that TDC improves previous Gaussian VAE discretization methods, such as TokenBridge (Wang et al., 2025).
Our contributions are summarized as follows:
  • (Section 3.1) We propose Gaussian Quant (GQ), a simple yet effective approach that constructs a VQ-VAE by training a constrained Gaussian VAE and converting it into a VQ-VAE without additional training.
  • (Section 3.2) We theoretically show that when the GQ codebook size matches the bits-back coding rate of the Gaussian VAE, the conversion error remains small.
  • (Section 3.3) We introduce the target divergence constraint (TDC) to train Gaussian VAEs for effective conversion, and show that GQ outperforms prior VQ-VAEs on both UNet and ViT architectures.
  • (Section 3.6) We show that TDC improves TokenBridge, an existing Gaussian VAE conversion methods.
@@ 2 Preliminaries
We denote the original image as $X$, the latent as $Z$, the encoder as $f(\cdot)$, and the decoder as $g(\cdot)$. We use $\log(\cdot)$ for the natural logarithm and $D_{KL}(\cdot||\cdot)$ for KL divergence. We denote the bits-back coding bitrate as $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$. We denote the base-2 logarithm as $\log_{2}(\cdot)$, with the corresponding bitrate $R_{i}^{\mathrm{bits}}=R_{i}/\log 2$.
@@ 2.1 Vector Quantized Variational Autoencoder
VQ-VAE (Van Den Oord et al., 2017) encodes images into integer tokens. For autoregressive generation, it uses deterministic encoding and a shared codebook. Specifically, VQ-VAE maintains a codebook $c_{1:K}$ of size $K$ with bitrate $\log K$. For each dimension $i$, the encoder output $f(x)_{i}$ is quantized to the closest codeword $c_{j}$ in $c_{1:K}$. Denoting distortion as $\Delta(\cdot,\cdot)$, the VQ-VAE objective is the rate-distortion function weighted by a Lagrange multiplier $\lambda$:
$\mathcal{L}_{VQ}=\lambda\underbrace{\log K}_{\textrm{bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(\hat{z}))]}_{\textrm{distortion}}+\mathcal{L}_{Reg},$
@@
where $c_{1:K}$ is the learned codebook, and $\mathcal{L}_{\mathrm{Reg}}$ is a regularization term that ensures VQ-VAE convergence, such as commitment loss, codebook loss (Van Den Oord et al., 2017), and Gumbel-Softmax loss (Sønderby et al., 2017).
2.2 Gaussian VAE and Bits-Back Coding
The Gaussian VAE is a special VAE (Kingma et al., 2013) with a prior $\mathcal{N}(0,I)$ and a fully factorized Gaussian posterior $q(Z|X)$. Its encoding simply involves sampling each latent dimension $z_{i}\sim q(Z_{i}|X)$. Assuming $\log p(X|Z=z)\propto(1/\lambda)\Delta(X,g(z))$, the negative evidence lower bound (ELBO) of the Gaussian VAE corresponds to a rate-distortion function, consisting of a bits-back coding bitrate term and a distortion term:
@@ 2.2 Gaussian VAE and Bits-Back Coding
The bits-back coding bitrate of $Z_{i}$ is defined as $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$ (Hinton and Van Camp, 1993), because when compressing $X$ losslessly, $z_{i}$ can be communicated using $R_{i}$ nats to arbitrary precision.
@@ 3 Gaussian Quant: Vector Quantization using Gaussian VAE
We propose an alternative approach to obtain a VQ-VAE: first train a constrained Gaussian VAE, then convert it into a VQ-VAE.
3.1 Converting a Gaussian VAE into a VQ-VAE
Given a Gaussian VAE, we convert it into a VQ-VAE by generating one-dimensional Gaussian noise as a codebook and quantizing the posterior mean independently for each latent dimension. The codebook is fixed once generated. Because it consists entirely of Gaussian samples, we call this approach Gaussian Quant (GQ). Specifically, we randomly generate a codebook $c_{1:K}\sim\mathcal{N}(0,1)$ shared across all dimensions. Then, for each dimension $i$, we select the codeword $c_{j}$ closest to the posterior mean $\mu_{i}$ and denote the quantized value as $\hat{z}_{i}$:
@@ 3.2 Theoretical Relationship between the Codebook Size and Quantization Error
Why GQ works and how to choose $K$ are not obvious. Theoretically, we show that GQ preserves the rate-distortion properties of the Gaussian VAE: when the codebook bitrate $\log K$ matches the bits-back coding rate $R_{i}$ of the Gaussian VAE, the quantization error is small. More specifically, we denote the event $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ as large quantization error. Then, we show that the probability of a large quantization error decays doubly exponentially as the codebook bitrate $\log K$ exceeds the bits-back coding rate $R_{i}$.
Theorem 3.1.
@@
Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i}$ and $\sigma_{i}$, respectively. We assume $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$. Given a fixed $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding rate. That is,
$\textup{when }\log K=R_{i}+t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\leq\exp{(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5})}.$ (4)
@@
Conversely, when the codebook bitrate $\log K$ is smaller than the bits-back coding rate $R_{i}$, the probability of a large quantization error increases exponentially toward $1$.
Theorem 3.2.
@@
Following Theorem 3.1, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ falls below the bits-back coding rate. That is,
$\textup{when }\log K=R_{i}-t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}.$ (5)
@@
Theorems 3.1 and 3.2 provide a principled guideline for choosing $K$: $\log K$ should be close to the bits-back rate $R_{i}$. In practice, setting $\log_{2}K=\left\lceil R_{i}^{\mathrm{bits}}\right\rfloor$ typically yields sufficiently small reconstruction error, where $\left\lceil\cdot\right\rfloor$ denotes rounding. Using a larger $K$ offers no additional benefit, while a smaller $K$ significantly increases the error.
@@
3.3 Training a Conversion-ready Gaussian VAE with Target Divergence Constraint
Now that we know how to convert a Gaussian VAE into a VQ-VAE, we discuss how to train a Gaussian VAE for effective conversion. To construct a VQ-VAE with a specific codebook size $K$, we need to train the Gaussian VAE so that $R_{i}$ is close to $\log K$ for all dimensions $i=1,\dots,d$.
To train a Gaussian VAE under this constraint, we propose a heuristic Target Divergence Constraint (TDC). Specifically, we set the target $R_{i}$ to $\log K$. For each dimension $i$, we apply a larger penalty if $R_{i}$ exceeds $\log K+\alpha$ and a smaller penalty if $R_{i}$ falls below $\log K-\alpha$, using different $\lambda$ values for each case. Here, $\alpha$ is a hyperparameter controlling the threshold:
$\mathcal{L}_{TDC}=\sum_{i=1}^{d}\lambda_{i}R_{i}+\mathbb{E}[\Delta(X,g(z))],$
$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&R_{i}<\log K-\alpha,\\ \lambda_{\textrm{mean}},&R_{i}\in[\log K\pm\alpha],\\ \lambda_{\textrm{max}},&R_{i}>\log K+\alpha.\end{array}\right$. (9)
To determine $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$, we initialize them to 1 and update them according to the relationship between the statistics of $R_{i}$ and $\log K\pm\alpha$ during each gradient descent step:
$\lambda_{\textrm{min}}=\beta\lambda_{\textrm{min}}\textrm{ if}\textrm{ min}_{i}\{R_{i}\}>\log K-\alpha\textrm{ else }\beta^{-1}\lambda_{\textrm{min}},$
$\lambda_{\textrm{mean}}=\beta\lambda_{\textrm{mean}}\textrm{ if}\textrm{ mean}_{i}\{R_{i}\}>\log K\textrm{ else }\beta^{-1}\lambda_{\textrm{mean}},$
$\lambda_{\textrm{max}}=\beta\lambda_{\textrm{max}}\textrm{ if}\textrm{ max}_{i}\{R_{i}\}>\log K+\alpha\textrm{ else }\beta^{-1}\lambda_{\textrm{max}},$
where $\beta$ is a hyperparameter controlling the update speed. To avoid numerical issues, we clip $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$ to the range $10^{-3},10^{3}]$ after each update.
3.4 Supporting Multi-dimensional Codebook
The vanilla VQ-VAE has three key parameters: codebook size $K$, codebook dimension $m$, and number of tokens $N$. Now GQ is able to construct VQ-VAE with codebook dimension $m=1$, we can extend GQ to codebook dimension $m>1$. Specifically, given a Gaussian VAE, we can quantize $m$ latents into a single large token with codebook size $\log_{2}K=\left\lceil\sum_{k=i}^{i+m}R_{k}^{\mathrm{bits}}\right\rfloor$. During quantization, we minimize the $\sigma_{i}$-weighted distance to find nearest neighbour, with the codebook $c_{1:K}$ sampled from $m$ dimensional factorized standard Gaussian:
@@ 3.4 Supporting Multi-dimensional Codebook
 $\textrm{ where }c_{1:K}\sim\mathcal{N}(0,I_{m}).$ (10)
However, for low-bitrate cases, this vanilla $m$-dimensional conversion leads to codebook collapse. This occurs because $|\mu_{i}|$ is bounded by $\sqrt{2R_{i}}$ (see Lemma A.1), so when $R_{i}$ is small, codebook vectors far from 0 are never selected. To address this, we introduce a regularization term weighted by $\omega$ that encourages the selection of $c_{j}$ farther from 0: